# Boundary control systems on a 1D spatial domainBouchra Elghazi, Birgit Jacob, and Hans Zwart

### v1

@@ 1 Introduction
The concept of well-posedness is fundamental in the study of boundary control and observation, as it forms the basis for further control and stability analysis. Over the years, the study of well-posedness of linear time-invariant systems has attracted considerable attention, leading to a significant amount of literature. Seminal contributions in this area include the works by Lasiecka and Triggiani [LasieckaTriggiani2000a, LasieckaTriggiani2000b] and by Staffans [Sta:05]. For further reading, we also refer to [MorrisCheng:99, MorrisCheng:03, CurtainWeiss1989, CurtZwa:20, JacZw:12, TucsnakWeiss:14, Weiss2001]. Informally speaking, well-posedness refers to the property that for every initial condition in the state space and any input function in a specified space of functions, the system has a unique state trajectory and a unique output function. Moreover, the output must belong to a specified space of functions, and both the state and the output must depend continuously on the initial state and on the input. Equivalently, it concerns the well-definedness and boundedness of the four mappings from initial state to input, initial state to final state, input to state, and input to output (also called transfer function). For boundary control and observation systems that are impedance passive, there exists a simple way of characterizing the well-posedness. In fact, if the system is internally well-posed, then the well-posedness is equivalent to the boundedness of the transfer function on a vertical line in the open right half plane. Notable examples within this class are port-Hamiltonian systems which provide a powerful framework for modelling as it views physical systems from an energy-based perspective [Schaft:06, SchJelt:14, SchMas:02].
@@
generates a $C_{0}$-semigroup on $X$, is equivalent to the well-posedness of the system, see [JacZw:12, ZwGorMasVill:10]. This class of systems covers the wave equation, the transport equation, and the Timoshenko beam equation. However, this result does not extend to the system (1), that models for example the Schrödinger equation or the Euler-Bernoulli beam equation, see the counterexample in [ElghJacZw:25]. In [ElghJacZw:25], the well-posedness of the system (1) was proven under the assumptions that $P_{1}=0$ and $\mathcal{H}$ is constant. In this paper, we extend the analysis to the general cas and we derive a simple verifiable conditions for well-posedness of the system (1).
@@ 3 Analysis of well-posedness
Let $s=r+i\omega\in\mathbb{C}$, $r>1$ and $u_{e,0}\in\mathbb{F}^{m}$. By $x_{0}$ we denote the solution of the ordinary differential equation
@@
The boundary control and observation system (16) is well-posed if the matrix
@@
If the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible, the well-posedness of the system (16) follows from the fact that we may introduce the additional input $v(t)=W_{B,2}\mathcal{H}\tau(x)$. The system with the extended input $\begin{bmatrix}u\\ v\end{bmatrix}$ is well-posed by Theorem 1 as $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible. Hence we may choose any input that is (locally) square integrable. In particular, we can choose $v\equiv 0$ and thus the boundary control and observation system (16) is well-posed if the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible.
@@ 4 Exact controllability and exact observability
[TucsnakWeiss:14, Definition 6.2.1 and Definition 6.2.12] The boundary control and observation system (3) is called exactly controllable (in finite time), if there exists a time $T>0$ such that for every $x_{1}\in X$ there exists a control function $u\in\mathrm{L^{2}}((0,T);\mathbb{F}^{m})$ such that the corresponding mild solution $x$ with $x_{0}=0$ satisfies
@@
$\mathcal{A}^{\ast}x$ $=\left(-P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}-P_{0}\right)\mathcal{H}x,$

### v3

@@ 1 Introduction
The concept of well-posedness is fundamental in the study of boundary control and observation, as it forms the basis for further control and stability analysis. Over the years, the study of well-posedness of linear time-invariant systems has attracted considerable attention, leading to a significant amount of literature. Seminal contributions in this area include the works by Lasiecka and Triggiani [LasieckaTriggiani2000a, LasieckaTriggiani2000b] and by Staffans [Sta:05]. For further reading, we also refer to [MorrisCheng:99, MorrisCheng:03, CurtainWeiss1989, CurtZwa:20, JacZw:12, TucsnakWeiss:14, Weiss2001]. Informally speaking, well-posedness refers to the property that for every initial condition in the state space and any input function in a specified space of functions, the system has a unique state trajectory and a unique output function. Moreover, the output must belong to a specified space of functions, and both the state and the output must depend continuously on the initial state and on the input. Equivalently, it concerns the well-definedness and boundedness of the four mappings from initial state to output, initial state to final state, input to state, and input to output (also called transfer function). For boundary control and observation systems that are impedance passive, there exists a simple way of characterizing the well-posedness. In fact, if the system is internally well-posed, then the well-posedness is equivalent to the boundedness of the transfer function on a vertical line in the open right half plane. Notable examples within this class are port-Hamiltonian systems which provide a powerful framework for modelling as it views physical systems from an energy-based perspective [Schaft:06, SchJelt:14, SchMas:02].
@@
generates a $C_{0}$-semigroup on $X$, is equivalent to the well-posedness of the system, see [JacZw:12, ZwGorMasVill:10]. This class of systems covers the wave equation, the transport equation, and the Timoshenko beam equation. However, this result does not extend to the system (1), that models for example the Schrödinger equation or the Euler-Bernoulli beam equation, see the counterexample in [ElghJacZw:25]. In [ElghJacZw:25], the well-posedness of the system (1) was proven under the assumptions that $P_{1}=0$ and $\mathcal{H}$ is constant. In this paper, we extend the analysis to the general case and we derive a simple verifiable conditions for well-posedness of the system (1).
@@ 3 Analysis of well-posedness
Let $s=r+i\omega\in\mathbb{C}$, $r>1$ and $u_{e,0}\in\mathbb{F}^{2n}$. By $x_{0}$ we denote the solution of the ordinary differential equation
@@
The boundary control and observation system (16) is well-posed and regular if the matrix
@@
If the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible, the well-posedness of the system (16) follows from the fact that we may introduce the additional input $v(t)=W_{B,2}\mathcal{H}\tau(x)$. The system with the extended input $\begin{bmatrix}u\\ v\end{bmatrix}$ is well-posed by Theorem 1 as $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible. Hence we may choose any input that is (locally) square integrable. In particular, we can choose $v\equiv 0$ and thus the boundary control and observation system (16) is well-posed if the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible. The system can be regular by extending the output and using the previous result.
Remark 3.7.
The matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible if and only if the matrix $\begin{bmatrix}W_{B,1}\\ W_{B,2}\end{bmatrix}\begin{bmatrix}0&0\\ I&0\\ 0&0\\ 0&I\end{bmatrix}$ is invertible, or equivantly, the matrix $\begin{bmatrix}Q_{12}&Q_{14}\\ Q_{22}&Q_{24}\end{bmatrix}$ is invertible, where
$\begin{bmatrix}W_{B,1}\\ W_{B,2}\end{bmatrix}=\begin{bmatrix}Q_{11}&Q_{12}&Q_{13}&Q_{14}\\ Q_{21}&Q_{22}&Q_{23}&Q_{24}\end{bmatrix}$.
Note that the square matrices $Q_{12}$, $Q_{14}$, $Q_{22}$, $Q_{24}$ are the matrices that act only on $(\mathcal{H}x)^{\prime}(1,t)$ and $(\mathcal{H}x)^{\prime}(0,t)$.
@@ 4 Exact controllability and exact observability
[CurtZwa:20, Definition 6.2.1 and Definition 6.2.12] The boundary control and observation system (3) is called exactly controllable (in finite time), if there exists a time $T>0$ such that for every $x_{1}\in X$ there exists a control function $u\in\mathrm{L^{2}}((0,T);\mathbb{F}^{m})$ such that the corresponding mild solution $x$ with $x_{0}=0$ satisfies
@@
$\mathcal{A}^{\ast}x$ $=\left(-P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}-P_{1}\frac{\partial}{\partial\zeta}-P_{0}\right)\mathcal{H}x,$